OpenStudy (anonymous):

Find the intercepts and then graph the line -4x+2y=8

OpenStudy (anonymous):

you have to set up 2 problems for this one

OpenStudy (anonymous):

x intercept = (-2,0) y itnercept = (0,4)

OpenStudy (anonymous):

To find your intercepts plug in zero. I change it to y=mx+b form first but that's my preference. y=2x+4

OpenStudy (anonymous):

0=2x+4, so x=-2. and y=2(0)+4 so y=4. Your intercepts will be (-2,0) and (0,4). Hope this helps.

OpenStudy (anonymous):

Thank you guys and what about the Graph

OpenStudy (anonymous):

Hold on I think I can do one in Fathom and post a picture.... I think

OpenStudy (anonymous):

Ok Samsterz can you also draw the graph

OpenStudy (anonymous):

Not sure if you can open this because I have a program for graphing on my computer, don't know if you can open it if you don't have it. It would be much easier to graph by hand though yes.

OpenStudy (anonymous):

ok sec

OpenStudy (anonymous):

there u go

OpenStudy (anonymous):

Thanks you guys bt How can i be a wizard like you .I am only good in English

OpenStudy (anonymous):

im good in english as well, honors english im taking

OpenStudy (anonymous):

To find the x-intercept, substitute in 0 for y and solve for x. 6x+5(0)=−30 Multiply 5 by each term inside the parentheses. More Detail Multiply 5 by the 0 inside the parentheses. 6x+5⋅0=−30 Multiply 5 by 0 to get 0. 6x+0=−30 6x+0=−30 Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression. 6x=−30 Divide each term in the equation by 6. 6x/6=−30/6 Simplify the left-hand side of the equation by canceling the common factors. x=−30/6 Simplify the equation. Cancel the common factor of 6 in −30/6. Reduce the expression −30/6 by removing a factor of 6 from the numerator and denominator. x=−5 x=−5 To find the y-intercept, substitute in 0 for x and solve for y. 6(0)+5y=−30 Multiply 6 by each term inside the parentheses. 0+5y=−30 Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression. 5y=−30 Divide each term in the equation by 5. Simplify the left-hand side of the equation by canceling the common factors. Cancel the common factor of 5 in 5y/5. Reduce the expression 5y/5 by removing a factor of 5 from the numerator and denominator. y=−305 y=−30/5 Solve the equation. Cancel the common factor of 5 in −30/5. Reduce the expression −305 by removing a factor of 5 from the numerator and denominator. y=−6 y=−6 These are the x and y intercepts of the equation 6x+5y=−30. x-intercept: (−5,0) y-intercept: (0,−6)